OpenStudy (anonymous):

perform the indicated operation of multiplication or division on the rational expressions and simplify: xy x ________ / _____________ x^2+11x+18 6x^2+7x-10

OpenStudy (anonymous):

invert the second bracket and multiply xy * 6x^2 + 7x + 10 ---------- -------------- x^2 + 11x + 18 x now check to see if the two quadratics can be factored

OpenStudy (anonymous):

6x^2 + 7x - 10 = (6x - 5)(x + 2) x^2 + 11x + 18 = (x + 9)(x + 2) sub these two factors in and cancel and your home (sorry i typed the top quadratic incorrectly in my last post i put + 10 instead of -10))